Hi guys, so I have the MW of the compound and it is 345 g/mol the compound is 0,2 M so it is 0,2 mol/liter and I need to calculate the mass?
1
Expert's answer
2021-08-24T06:16:34-0400

"Molarity =\frac{Moles}{Volume}"


"Volume=\frac{ moles}{Molarity}"


"Volume=\frac{0.2}{0.2}=1liter"


"Mass= 0.2\u00d7345= 69grams"